# Better by Comparison: Retrieval-Augmented Contrastive Reasoning for Automatic Prompt Optimization

### v1

@@ Abstract
Automatic prompt optimization has recently emerged as a strategy for improving the quality of prompts used in Large Language Models (LLMs), with the goal of generating more accurate and useful responses. However, most prior work focuses on direct prompt refinement or model fine-tuning, overlooking the potential of leveraging LLMs’ inherent reasoning capability to learn from contrasting examples. In this paper, we present Contrastive Reasoning Prompt Optimization (CRPO), a novel framework that formulates prompt optimization as a retrieval-augmented reasoning process. Our approach retrieves top-$k$ reference prompts from the HelpSteer2 dataset, an open-source collection annotated for helpfulness, correctness, coherence, complexity, and verbosity, and constructs two complementary optimization paradigms: (1) tiered contrastive reasoning, where the LLM compares high-, medium-, and low-quality prompts to refine its own generation through reflective reasoning, and (2) multi-metric contrastive reasoning, where the LLM analyzes the best prompts along each evaluation dimension and integrates their strengths into an optimized prompt. By explicitly contrasting high- and low-quality exemplars, CRPO enables the model to deduce why certain prompts succeed while others fail, thereby achieving more robust and interpretable optimization. Experimental results on the HelpSteer2 benchmark demonstrate that CRPO significantly outperforms baselines. Our findings highlight the promise of contrastive, retrieval-augmented reasoning for advancing automatic prompt optimization.
Index Terms:
Automatic Prompt Optimization, Contrastive Reasoning, Retrieval-Augmented Generation, Large Language Models (LLMs), Helpfulness Alignment.
@@ I Introduction
Early approaches to automatic prompt optimization explored continuous soft prompt tuning [3, 4, 5] or discrete search over token combinations [6, 7, 8]. While effective in scenarios with access to model gradients (white-box settings), these methods are inapplicable to black-box LLMs (e.g., those accessed via API), where internal states are inaccessible. More recently, optimization has shifted toward black-box prompting methods, where models iteratively explore candidate prompts through trial-and-error [9, 10]. Representative frameworks include PromptAgent [11], which formulates prompt optimization as a Monte Carlo Tree Search planning problem to simulate human trial-and-error exploration; Hierarchical Multi-Agent Workflows (HMAW) [12], which adopt a CEO–Manager–Worker structure to decompose optimization into hierarchical roles; and MASS [13], which leverages topology optimization across multiple agents for selecting effective prompts.
Another line of work views LLMs themselves as optimizers. For instance, OPRO [14] treats prompt optimization as a natural language optimization task, where the LLM iteratively generates new solutions conditioned on previously evaluated prompts. Similarly, PromptBreeder [15] evolves a population of prompts through mutation and selection to improve robustness, and AMPO [16] applies a multi-branch structure that grows and prunes prompts like a tree.
@@
To address these limitations, we introduce Contrastive Reasoning Prompt Optimization (CRPO), a retrieval-augmented framework that explicitly leverages contrastive reasoning across prompts of varying quality. Instead of fine-tuning the LLM, our method retrieves top-$k$ reference prompts from the HelpSteer2 dataset [17], which is an open-source benchmark annotated for helpfulness, correctness, coherence, complexity, and verbosity, and performs two novel optimization strategies: (1) tiered contrastive reasoning, where the model reflects on high-, medium-, and low-quality exemplars to refine its own generation, and (2) multi-metric contrastive reasoning, where the best prompts along individual metrics are analyzed and integrated into an optimized prompt. CPRO enhances prompt quality without requiring model parameter updates by reasoning explicitly about why certain prompts succeed and others fail.
Extensive experiments on HelpSteer2 show that CPO consistently outperforms our baselines, including (1) direct generation, (2) Chain-of-Thought prompting (CoT) [18], and simple retrieval-based augmentation. [^1] Detailed implementations of baselines are detailed in Section IV. These results demonstrate the promise of contrastive, retrieval-augmented optimization for aligning LLM outputs toward more helpful, factually correct, and coherent responses.
@@ II Dataset
We conduct our experiments on the HelpSteer2 dataset [17], a benchmark designed for evaluating prompt optimization and response helpfulness in Large Language Models (LLMs). The dataset consists of human-annotated prompt–response pairs and is divided into training and validation splits. Specifically, it contains a training set of 20.3k rows and a validation set of 1.04k rows, which we adopt for retrieval and evaluation, respectively.
@@ III Methodology
We formulate the problem of prompt optimization problem not as directly fine-tuning model parameters, but as a task that maximizes the inherent reasoning capability of LLMs by facilitating learning from reference examplars. Specifically, our framework Contrastive Reasoning Prompt Optimization (CRPO) first retrieves the top-$k$ relevant prompts from the HelpSteer2 training set, and then applies contrastive reasoning to construct optimized prompts. We design two complementary variants, illustrated in Figure 1.
III-A Retrieval of Reference Prompts
Given an input query $q$, CRPO retrieves a set of reference prompts $\{p_{1},\dots,p_{k}\}$ using a retriever:
$\mathcal{R}(q)=\{p_{1},\dots,p_{k}\},\quad p_{i}\in\text{HelpSteer2}.$ (1)
Each $p_{i}$ is annotated along five evaluation dimensions $\mathcal{M}=\{\text{help},\text{corr},\text{coh},\text{comp},\text{verb}\}$. To ensure sufficient coverage across all dimensions, we require $k\geq 5$, so that at least one candidate prompt is available per metric. These exemplars serve as contrasting cases that enable the LLM to perform explicit reasoning.
@@ III-B Variant 1: Tiered Contrastive Reasoning
We partition $\mathcal{R}(q)$ into three tiers according to overall quality scores. Specifically, for each retrieved prompt $p_{i}$, we compute its average score across all five evaluation dimensions:
$\text{Avg}(p_{i})=\frac{1}{|\mathcal{M}|}\sum_{m\in\mathcal{M}}\text{Score}(p_{i},m),$ (2)
where $\mathcal{M}=\{\text{help},\text{corr},\text{coh},\text{comp},\text{verb}\}$. Prompts are then partitioned into high-quality ($P^{H}$), medium-quality ($P^{M}$), and low-quality ($P^{L}$) tiers based on quantile thresholds of $\text{Avg}(p_{i})$. The optimized prompt $p^{*}$ is generated via contrastive reasoning:
@@ III-C Variant 2: Multi-Metric Contrastive Reasoning
For each metric $m\in\mathcal{M}$, we select the top prompt:
$P^{m}=\arg\max_{p_{i}\in\mathcal{R}(q)}\text{Score}(p_{i},m).$
@@
where $\text{Integrate}(\cdot)$ encourages the LLM to combine complementary strengths across evaluation axes.
III-D Evaluation Method
Effectiveness is assessed by directly feeding the original prompt $p$ and the optimized prompt $p^{*}$ into the evaluation model $\mathcal{E}$ as a pair. The evaluation model then outputs comparative scores across the five annotated dimensions $\mathcal{M}$:
$\text{Score}(p,p^{*})=\big{[}s_{\text{help}},s_{\text{corr}},s_{\text{coh}},s_{\text{comp}},s_{\text{verb}}\big{]},$ (5)
where each $s_{m}\in[0,4]$ for $m\in\mathcal{M}$.
To ensure consistency, these raw scores are normalized to the $[0,1]$ range and aggregated as described in Section 4. Improvements are thus attributed to contrastive reasoning in CRPO, since the evaluation model judges $p^{*}$ relative to $p$ under identical conditions.
Algorithm 1 Contrastive Reasoning Prompt Optimization (CRPO)
1:Query $q$, Retriever $\mathcal{R}$, LLM $f_{\theta}$, Evaluation Model $\mathcal{E}$, Metrics $\mathcal{M}$, Top-$k$
2:// Retrieval Stage
3:Retrieve prompts: $\mathcal{R}(q)=\{p_{1},\dots,p_{k}\}$
4:// Variant 1: Tiered Contrastive Reasoning
5:Partition $\mathcal{R}(q)$ into $P^{H},P^{M},P^{L}$
[…]
7:// Variant 2: Multi-Metric Contrastive Reasoning
8:for each metric $m\in\mathcal{M}$ do
9:  $P^{m}\leftarrow\arg\max_{p_{i}\in\mathcal{R}(q)}\text{Score}(p_{i},m)$
10:end for
[…]
12:// Evaluation
13:Evaluate: $\text{Score}(\hat{y})=\mathcal{E}(p,p^{*})$
@@ IV-A Setup
We conduct experiments on the HelpSteer2 dataset to evaluate the effectiveness of our proposed Contrastive Reasoning Prompt Optimization (CRPO). All experiments are designed to ensure fairness in comparison with baseline methods by using the same retrieval pool and evaluation settings.
@@ IV-A1 LLM and Retrieval Models
For generations, we utilize two representative large language models:
  • GPT-4o [19]: accessed via API, configured with temperature =0=0 to ensure deterministic outputs. All other hyperparameters are kept at their default settings.
  • LLaMA 3-8B [20]: an open-source instruction-tuned model, also set with temperature =0=0 for reproducibility.
For retrieval, we adopt the BM25 retriever, which selects top-$k$ reference prompts $\{p_{1},\dots,p_{k}\}$ for each query. We set $k=10$ across all experiments, ensuring sufficient coverage across the $5$ evaluation dimensions while maintaining efficiency.
GPT-4o
Direct Generation 0.3655 0.4349 0.7669 0.4050 0.6638 0.5272
Chain-of-Thought (CoT) 0.3660 0.4329 0.7675 0.3998 0.6666 0.5266
CRPO-Tiered Contrastive Reasoning 0.5253 0.6065 0.8817 0.4474 0.7168 0.6355
CRPO-Multi-Metric Contrastive Reasoning 0.5158 0.5964 0.8758 0.4316 0.6780 0.6195
LLaMA 3-8B
Direct Generation 0.3438 0.4101 0.7526 0.4025 0.6602 0.5138
Chain-of-Thought (CoT) 0.3432 0.4067 0.7554 0.4107 0.6657 0.5163
Retrieval Augmented Generation (RAG) 0.3422 0.4161 0.7607 0.4003 0.6492 0.5137
CRPO-Tiered Contrastive Reasoning 0.4224 0.5023 0.8092 0.4243 0.6685 0.5654
CRPO-Multi-Metric Contrastive Reasoning 0.3990 0.4711 0.7989 0.4172 0.6564 0.5485
TABLE II: Comparison across five evaluation metrics—helpfulness, correctness, coherence, complexity, and verbosity—and their mean (Avg. Score). Results are reported separately for GPT-4o and LLaMA-3-8B. The best value within each language model is bold, the second best is underlined, and methods marked with denote our proposed CRPO variants.
IV-A2 Evaluation Model
For evaluation, we adopt the ArmoRM-Llama3-8B-v0.1 reward model [21], a recently proposed interpretable multi-objective reward model that leverages absolute ratings across human-interpretable dimensions. ArmoRM has demonstrated superior performance on the benchmark dataset, achieving over $90\%$ accuracy and ranking among the top reward models available. We adopt ArmoRM as the automatic judge for assessing the quality of generated responses.
IV-A3 Evaluation Metrics
Each response in the HelpSteer2 dataset is annotated across $5$ dimensions, with scores ranging from 0 to $4$. To normalize across dimensions, we first scale each score by dividing by $4$, mapping it to the range $[0,1]$.
Formally, for a given response $\hat{y}$, the normalized score is computed as:
$\text{Score}(\hat{y})=\frac{1}{|\mathcal{M}|}\sum_{m\in\mathcal{M}}\frac{s_{m}(\hat{y})}{4},$ (6)
[…]
IV-A4 Baselines
We compare CRPO against three representative baselines:
  • Direct Generation: the LLM generates a response directly from the query without any retrieval or optimization.
  • Chain-of-Thought (CoT): the LLM is instructed to reason step-by-step before producing a final response.
  • Retrieval-Augmented Generation (RAG): the LLM receives the query along with the top-$k$ retrieved prompts (with $k=10$, for fair comparison).
To ensure fairness, all methods are evaluated using the same retrieval pool and judged by the same evaluation model. The quality of generated responses is mapped to the metrics provided in HelpSteer2, which includes five dimensions: helpfulness, correctness, coherence, complexity, and verbosity.
@@ IV-B Main Experiment Results
Baseline methods demonstrate modest, albeit limited improvements. Direct generation often produces shallow or inconsistent outputs, since no external context is used. Chain-of-Thought (CoT) prompting encourages stepwise reasoning but tends toward verbosity and repetition without improving factuality. Retrieval-Augmented Generation (RAG) provides contextual grounding through exemplars, but introduces redundancy and lacks a clear mechanism to filter high-quality signals. Thus, while baselines improve over the previous ones, they remain insufficient for producing balanced prompts.
In contrast, CRPO addresses these shortcomings through explicit contrastive reasoning. The multi-metric variant integrates the strongest exemplars along helpfulness, correctness, coherence, complexity, and verbosity, ensuring complementary strengths are preserved. The tiered variant contrasts high-, medium-, and low-quality prompts, adopting strengths from the best, avoiding weaknesses from the worst, and stabilizing with medium-quality anchors. Together, these strategies yield prompts that are more robust, interpretable, and human-aligned than those produced by direct generation, CoT, or RAG.
@@ IV-C1 Trimaximal-Prompt Selection (TPS)
To assess the role of contrastive reasoning, we conduct an ablation where CRPO is reduced to using only the top-$3$ highest-ranked prompts, without tiered or multi-metric reasoning. As shown in Figure 2, this simplified method performs worse across evaluation metrics, confirming that simple ranking is insufficient.
In contrast, both CRPO variants consistently outperform the ablated setting by explicitly reasoning over contrasts—adopting strengths from high-quality prompts, avoiding weaknesses from low-quality ones, and integrating complementary aspects across dimensions. This result shows that the gains of CRPO stem not from retrieval alone, but from its reflective reasoning process, which yields more stable and interpretable optimization.

### v2

@@ Abstract
Automatic prompt optimization has recently emerged as a strategy for improving the quality of prompts used in Large Language Models (LLMs), with the goal of generating more accurate and useful responses. However, most prior work focuses on direct prompt refinement or model fine-tuning, overlooking the potential of leveraging LLMs’ inherent reasoning capability to learn from contrasting examples. In this paper, we present Contrastive Reasoning Prompt Optimization (CRPO), a novel framework that formulates prompt optimization as a retrieval-augmented reasoning process. Our approach retrieves top-$k$ reference prompt-response pairs from the HelpSteer2 dataset, an open-source collection where each response is annotated for helpfulness, correctness, coherence, complexity, and verbosity, and constructs two complementary optimization paradigms: (1) tiered contrastive reasoning, where the LLM compares high-, medium-, and low-quality exemplars (both prompts and responses) to refine its own generation through reflective reasoning, and (2) multi-metric contrastive reasoning, where the LLM analyzes the best exemplars along each evaluation dimension and integrates their strengths into an optimized prompt. By explicitly contrasting high- and low-quality exemplars, CRPO enables the model to deduce why certain prompts succeed while others fail, thereby achieving more robust and interpretable optimization. Experimental results on the HelpSteer2 benchmark demonstrate that CRPO significantly outperforms baselines. Our findings highlight the promise of contrastive, retrieval-augmented reasoning for advancing automatic prompt optimization.
@@ I Introduction
Early prompt optimization methods focused on continuous soft prompt tuning [3, 4, 5] or discrete token search [6, 7, 8], which were effective in white-box settings but inapplicable to black-box LLMs accessed via API. More recent work has shifted to black-box approaches that explore prompts through trial-and-error, such as PromptAgent’s Monte Carlo Tree Search [9], hierarchical multi-agent workflows [10], and multi-agent topology optimization [11]. In parallel, some methods treat LLMs themselves as optimizers, for instance, OPRO [12] frames optimization as iterative natural language refinement, while evolutionary approaches like PromptBreeder [13] and branching structures such as AMPO [14] adapt prompts through mutation, selection, and pruning.
@@
To address these limitations, we introduce Contrastive Reasoning Prompt Optimization (CRPO), a retrieval-augmented framework that explicitly leverages contrastive reasoning across prompts of varying quality. Unlike methods that only consider prompt text and scores, CRPO retrieves top-$k$ reference prompt-response pairs from the HelpSteer2 dataset [15] using sparse retrieval over prompt text, where each pair is annotated for helpfulness, correctness, coherence, complexity, and verbosity. By incorporating both the retrieved prompts and their responses, CRPO allows the model to reason not only about how prompts are phrased but also how they shape the quality of generated outputs. It then performs two novel optimization strategies: (1) tiered contrastive reasoning, where the model reflects on high-, medium-, and low-quality exemplars to refine its own generation, and (2) multi-metric contrastive reasoning, where the best exemplars along individual metrics are analyzed and integrated into an optimized prompt.
CRPO improves prompt quality without model updates by explicitly reasoning about why prompts succeed or fail; on HelpSteer2 it consistently outperforms direct generation, Chain-of-Thought prompting (CoT) [16], and simple retrieval augmentation [^1] Detailed implementations of baselines are detailed in Section IV., underscoring the promise of contrastive, retrieval-augmented optimization for aligning LLM outputs toward more helpful, factual, and coherent responses.
@@ II Dataset
We conduct our experiments on the HelpSteer2 dataset [15], a benchmark designed for evaluating prompt optimization and response helpfulness in Large Language Models (LLMs). The dataset consists of human-annotated prompt-response pairs and is divided into training and validation splits. Specifically, it contains a training set of 20.3k rows and a validation set of 1.04k rows, which we adopt for retrieval and evaluation, respectively.
@@
GPT-4o
Direct Generation 0.3616 0.4700 0.7723 0.3280 0.4913 0.4846
Chain-of-Thought (CoT) 0.2981 0.3958 0.7169 0.2888 0.4709 0.4341
CRPO-Tiered Contrastive Reasoning 0.5274 0.6006 0.8711 0.4360 0.6506 0.6171
CRPO-Multi-Metric Contrastive Reasoning 0.5219 0.6274 0.8876 0.4386 0.6982 0.6347
LLaMA-3-8B
Direct Generation 0.2616 0.3636 0.7078 0.2942 0.4421 0.4139
Chain-of-Thought (CoT) 0.1600 0.2545 0.6291 0.2421 0.3812 0.3334
Retrieval Augmented Generation (RAG) 0.3990 0.4711 0.7989 0.3722 0.5814 0.5245
CRPO-Tiered Contrastive Reasoning 0.4224 0.5023 0.8092 0.3993 0.6435 0.5554
CRPO-Multi-Metric Contrastive Reasoning 0.4264 0.5070 0.8159 0.4146 0.6404 0.5609
TABLE II: Comparison across five evaluation metrics—helpfulness, correctness, coherence, complexity, and verbosity—and their mean (Avg. Score). Results are reported separately for GPT-4o and LLaMA-3-8B. The best value within each language model is bold, the second best is underlined, and methods marked with † denote our proposed CRPO variants.
@@ III Methodology
We formulate the prompt optimization problem not as directly fine-tuning model parameters, but as a task that maximizes the inherent reasoning capability of LLMs by facilitating learning from reference exemplars. Specifically, our framework Contrastive Reasoning Prompt Optimization (CRPO) first retrieves the top-$k$ relevant prompt-response pairs from the HelpSteer2 training set, and then applies contrastive reasoning to construct optimized prompts. We design two complementary variants, illustrated in Figure 1.
III-A Retrieval of Reference Prompt–Response Pairs
Given an input query $q$, CRPO retrieves a set of reference prompt–response pairs
$\mathcal{R}(q)=\{(p_{1},r_{1}),\dots,(p_{k},r_{k})\},\quad(p_{i},r_{i})\in\text{HelpSteer2},$
where $p_{i}$ denotes a prompt and $r_{i}$ its associated response. Each pair is annotated along five evaluation dimensions $\mathcal{M}=\{\text{help},\text{corr},\text{coh},\text{comp},\text{verb}\}$. To ensure sufficient coverage across all dimensions, we require $k\geq 5$, so that at least one candidate pair is available per metric. These exemplars (including both prompts and responses) serve as contrasting cases that enable the LLM to perform explicit reasoning.
@@ III-B Variant 1: Tiered Contrastive Reasoning
We partition $\mathcal{R}(q)$ into 3 tiers according to overall quality scores. Specifically, for each retrieved pair $(p_{i},r_{i})$, we compute its average score across all five evaluation dimensions:
$\text{Avg}(p_{i},r_{i})=\frac{1}{|\mathcal{M}|}\sum_{m\in\mathcal{M}}\text{Score}((p_{i},r_{i}),m),$ (1)
where $\mathcal{M}=\{\text{help},\text{corr},\text{coh},\text{comp},\text{verb}\}$. Pairs are then partitioned into high-quality ($P^{H}$), medium-quality ($P^{M}$), and low-quality ($P^{L}$) tiers based on quantile thresholds of $\text{Avg}(p_{i},r_{i})$. The optimized prompt $p^{*}$ is generated via contrastive reasoning:
@@ III-C Variant 2: Multi-Metric Contrastive Reasoning
For each metric $m\in\mathcal{M}$, we select the top pair:
$P^{m}=\arg\max_{(p_{i},r_{i})\in\mathcal{R}(q)}\text{Score}((p_{i},r_{i}),m).$
@@
where $\text{Integrate}(\cdot)$ encourages the LLM to combine complementary strengths across evaluation axes, grounded in both prompts and their associated responses.
@@ IV-A Setup
We evaluate CRPO on the HelpSteer2 dataset, ensuring fairness by using the same retrieval pool and evaluation settings as the baselines.
@@ IV-A1 LLM and Retrieval Models
We use GPT-4o and LLaMA-3-8B, both with temperature 0 and otherwise default hyperparameters. BM25 is specifically used for prompt retrieval, selecting top-$k$ reference prompts $\{p_{1},\dots,p_{k}\}$ per query with $k=10$ to balance coverage across the 5 evaluation dimensions and efficiency.
IV-A2 Evaluation Model and Metrics
We employ the ArmoRM-Llama3-8B-v0.1 reward model as an interpretable multi-objective judge ($90\%$ benchmark accuracy), which automatically scores response quality.
For evaluation metrics, we assess effectiveness by feeding the optimized prompt $p^{*}$ together with its generated response $r^{*}$ to the evaluation model $\mathcal{E}$, which returns five HelpSteer2 scores—helpfulness, correctness, coherence, complexity, verbosity—each in $[0,4]$:
$\text{Score}(p^{*},r^{*})=\big[s_{\text{help}},s_{\text{corr}},s_{\text{coh}},s_{\text{comp}},s_{\text{verb}}\big].$ (4)
Each score is normalized by dividing by 4 and then averaged to yield a final $[0,1]$ quality score:
$\mathrm{Score}(p^{*},r^{*})=\tfrac{1}{|\mathcal{M}|}\sum_{m\in\mathcal{M}}\tfrac{s_{m}(p^{*},r^{*})}{4}.$ (5)
$(p^{*},r^{*})$ is judged under identical conditions to the baselines.
IV-A3 Baselines
We compare our framework against three representative baselines, all of which optimize the prompt based only on the prompt text and evaluation scores, without incorporating structured reasoning assets:
  • Direct Generation: The LLM directly produces an optimized prompt from the input query, without retrieval or iterative refinement.
  • Chain-of-Thought (CoT): The LLM is instructed to perform step-by-step reasoning on the input query before generating an optimized prompt, relying solely on the original prompt text and scores.
  • Retrieval-Augmented Generation (RAG): The LLM receives the input query along with the top-$k$ retrieved training prompts (with $k=10$ for fair comparison) and generates an optimized prompt based on these examples and their scores, but without leveraging the associated reasoning assets.
@@ IV-B Main Experiment Results
Baseline methods demonstrate modest, albeit limited improvements. Direct Generation often produces shallow or inconsistent outputs, since no external context is used. Chain-of-Thought (CoT) prompting encourages stepwise reasoning but tends toward verbosity and repetition without improving factuality. Retrieval-Augmented Generation (RAG) provides contextual grounding through retrieved prompts, but introduces redundancy and lacks a clear mechanism to filter high-quality signals. Importantly, all three baselines optimize prompts only with respect to prompt text and evaluation scores, without leveraging the responses associated with those prompts. As a result, while baselines improve over one another incrementally, they remain insufficient for producing balanced and robust outputs.
In contrast, CRPO addresses these shortcomings through explicit contrastive reasoning over prompt–response pairs. The multi-metric variant integrates the strongest exemplars along helpfulness, correctness, coherence, complexity, and verbosity, ensuring complementary strengths are preserved. The tiered variant contrasts high-, medium-, and low-quality pairs, adopting strengths from the best, avoiding weaknesses from the worst, and stabilizing with medium-quality anchors. Together, these strategies yield optimized prompts—and consequently generated responses—that are more robust, interpretable, and human-aligned than those produced by Direct Generation, CoT, or RAG.
@@ IV-C1 Trimaximal-Prompt Selection (TPS)
To assess the role of contrastive reasoning, we conduct an ablation where CRPO is reduced to using only the top-$3$ highest-ranked prompts, as shown in Figure 2. In contrast, both CRPO variants consistently outperform the ablated setting by explicitly reasoning over contrasts—adopting strengths from high-quality prompts, avoiding weaknesses from low-quality ones, and integrating complementary aspects across dimensions. This result shows that the gains of CRPO stem not from retrieval alone, but from its reflective reasoning process, which yields more stable and interpretable optimization.